Determining A simple interest rate
you borrow $2500 froma friend and promise to pay back $2655 in six months. What simple interest rate will you pay?
example:
2655=2500[1 + r(0.5)]
2655=2500 + 1250r
I don't know how they got 1250r (distributive property)
Please explain
Thank you

Answer by rmromero(383)

you borrow $2500 froma friend and promise to pay back $2655 in six months.  What simple interest rate will you pay?

What is asked in the problem?
  Simple interest 

Given :
   money borrowed $ 2500
   Pay back is $2655

Representation
  P = the money borrowed $2500
  I = interest
  R = rate
  t = time
  A = pay back

Equation:
    A = P (1 + rt)
 2655 = 2500[1 + r(0.5)]
 2655 = 2500 + 1250r

You lack this part
 2655 - 2500 = 1250r
         155 = 1250r    divide both sides by 1250
    155/1250 = 1250r/1250
       0.124 = r

To change decimal to percent 
                0.124
    0.124  =   ______     = 12.4%  --> answer

                 100
check your answer if it is correct



Rachel